OpenStudy (anonymous):

A discriminant of 0 indicates what type of roots?

zepdrix (zepdrix):

Ummm, I suppose it means that you only have 1 root, and it's a real root. Example:\[\large 0=x^2-2x+1\]Throwing this into the Quadratic Formula gives us,\[\large x=\frac{2 \pm \sqrt{2^2-4(1)(1)}}{2} \qquad \rightarrow \qquad x=1\]

OpenStudy (anonymous):

One Rational Double

zepdrix (zepdrix):

Yes, that sounds right :) It's a repeated, or "double", root since there is only one.